### v3

@@ 3 Approach
Figure 2: The overall framework of SAR-KGPA. The proposed Knowledge-Guided Predictive Architecture for SAR target recognition (SAR-KGPA) uses locally masked patches to predict the multi-scale SAR feature representations of unseen patches. The MIM structure uses the Vision Transformer (ViT) in MAE. The DNN encoder extracts deep features of masked patches, and the DNN predictor predicts SAR features of unseen patches. The target encoder uses the gradient-by-ratio method to map SAR images from pixel space to feature space, thus extracting the target shapes and avoiding speckle noise interference in SAR images. Local patches and multi-scale features are designed for multi-scale small targets in remote sensing.
The proposed Knowledge-Guided Predictive Architecture for SAR ATR (SAR-KGPA) is illustrated in Figure 2. Our overall objective is as follows: given local masked patches, predict feature representations in the SAR domain. Therefore, the model can achieve SAR image interpretation by recognizing the contextual relationship between the target parts and the surrounding area.
As shown in Figure 2, the input is the single-channel SAR magnitude image since most target datasets are magnitude pictures. We randomly select different local patches and add mask tokens according to LoMaR (Chen et al., 2022a). Vision Transformers (ViT) (Dosovitskiy et al., 2020) in MAE are used as the DNN-encoder and DNN-predictor. Each local masked patch is fed into the MIM structure to predict the target feature of mask tokens.
Importantly, the special SAR coherent imaging mechanism requires domain knowledge to predict target features. SAR coherent imaging results in multiplicative scattering noise in the magnitude image, and we use the gradient-by-ratio (Dellinger et al., 2014; Song et al., 2016) to suppress the speckle noise and extract the targets’ edge and shape. Moreover, we employ a multi-scale setting for various targets in the pre-training set.
@@ 3.1 Local Masked Patches
We aim to learn the contextual relationship of various targets. Whereas such relationships normally remain in small areas due to the large scale and small target in remote sensing imagery. Therefore, we perform masking in the local patches following LoMaR. Although LoMaR is designed to improve computational efficiency, we find this local setting is suited for small targets in remote sensing. According to the LoMaR, mask tokens and relative positional encoding (Wu et al., 2021) are added to the encoder input and MAE structure to enhance the feature representation.
@@ 3.2 Target Encoder
Due to the SAR imaging mechanism (Moreira et al., 2013), the images’ amplitude values contain multiplicative speckle noise. The speckle noise is related to the sensor resolution and is caused by the coherent superposition of different scatterers within the scattering cell. Therefore, reconstructing image pixels for SAR images receives noise interference and fails to learn a high-quality feature representation. This problem of speckle noise can be solved by filtering and feature extraction. However, filtering has a conflict between noise reduction and image details, and we use feature extraction. Some studies have proposed different types of target features in computer vision, mainly divided into traditional manual features (Wei et al., 2022) and deep learning features (Assran et al., 2023). However, these methods are difficult to migrate to SAR target recognition directly, so we combine the traditional local description feature extraction methods (Dellinger et al., 2014; Song et al., 2016; Dong et al., 2020) in the SAR domain.
Gradient computation. The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in the SAR image, and previous studies (Touzi et al., 1988; Bovik, 1988) have shown that the computing ratio is more suitable for multiplicative noise. Here, we use the simplest ratio of average (ROA) (Touzi et al., 1988) not to blur image details:
@@
where $R_{i}$ denotes the average ratio at different directions and $M_{1}(i)$ and $M_{2}(i)$ are the area averages on opposite sides of the current pixel along direction $i$. $i=1$ is the horizontal direction and $i=3$ means the vertical direction. As shown in Figure 2, the area averages can be computed from the input image and four fixed convolution kernels.
The use of logarithms can solve the vertical gradient calculation (Dellinger et al., 2014), and the horizontal and vertical gradients are defined as:
@@
where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient. The gradient magnitude is $G_{\text{m}}=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}}$.
Multi-Scale Feature. SAR-HOG (Song et al., 2016) has discussed the kernel sizes for SAR vehicle target recognition and given the best single size for the MSTAR vehicle dataset. However, a single-size kernel cannot accommodate various targets when the dataset is extended to different targets and scenes. Due to the dynamic range of contextual information for various targets in remote sensing (Li et al., 2023c), a multi-scale feature is constructed with convolutional kernels of different sizes in Figure 1. We set the kernel scale $r$ equal to 5, 7, 13, and 17, and the whole convolutional kernel size is odd square $12r+$ 1. As shown in Figure 1, small-scale convolutional kernels provide a finer extraction of small target contours, while large-scale convolutional kernels focus on large targets and scene edges and noise suppression.
@@ 3.3 Implementation
Given SAR images, we randomly sample several square windows at different spatial locations as local patches and mask each patch with mask tokens at a fixed percentage. Then, each local masked patch’s visible and unseen parts are fed into the MIM structure, whose encoder and predictor are applied with learnable relative positional encoding in the self-attention layer. Meanwhile, the whole SAR image is provided to the target encoder. The average values in four directions are obtained after multi-scale convolution in four directions. Then, the gradients by ratio are obtained as multi-scale SAR features by the ratio and logarithm operations. The loss function computes the mean squared error in the feature space between the DNN features and SAR multi-scale features for masked patches. In addition to the major improvement of the target encoder, there are two other minor.
@@
Decoder design. While some work (Chen et al., 2022a; Wei et al., 2022) has shown that it is feasible to use a simple linear head as a decoder or predictor for self-supervised learning in RGB images, we find that a ViT with eight layers in the MAE is less disturbed by SAR image noise than a linear layer, allowing the encoder to extract deep semantic features of targets.
@@ 4 Experiments
This section analyzed the proposed architecture. We first described the pre-training and few-shot classification experimental setting in Sec. 4.1. Then, the proposed SAR-KGPA was compared with other methods in Sec. 4.2. We then discussed the local mask region in Sec. 4.3 and performed an ablation study for the SAR-KGPA with different target features in Sec. 4.4. In the end, Sec. 4.5 discussed the scaling experiment of our methods and showed improved performance with increasing data volume.
@@ 4.1 Dataset and Experimental Settings
Table 1: Description of SAR datasets used for pre-training. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Large SAR imagery in the detection contains more target and scene types than the annotation.
@@
Figure 3: Datasets for the pre-training contain different targets, scenes, and sensors. MSAR is the satellite-based dataset of ground and sea targets; SAR-Ship-Dataset is the satellite-based dataset of sea targets; SARSim is the multi-angle simulation dataset of vehicle targets; and SAMPLE is the simulated and measured dataset of vehicle targets.
Pre-training. Pre-training datasets are designed to ensure rich information about targets, scenes, and sensors. Therefore, we chose four datasets covering popular targets (vehicles, ships, aircraft, etc), complex backgrounds (ground and ocean), and various sensors (satellite-borne, airplane-borne, and simulation) in Table 1 and Figure  3.
  1. $\bullet$
@@ 3. $\bullet$
Experimental settings. Table 2 gives the default setting. The pre-training is applied on 4 NVIDIA V100 GPUs with 200 epochs and 300 batch sizes, and the backbone model is ViT-Base. Other hyperparameters of each method use the default settings from their papers and codes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness. All models use the same training settings in downstream classification tasks. The few-shot learning setting is based on the Dassl toolbox (Zhou et al., 2023a) and is averaged over 10 random experiments.

### v4

@@ 3 Approach
Figure 2: The overall framework of SAR-JEPA. (a) In the pre-training stage, our Joint-Embedding Predictive Architecture for SAR target recognition (SAR-JEPA) uses local masked patches to predict the multi-scale SAR feature representations $f_{\rm SAR}$ of unseen patches. The MIM structure uses the Vision Transformer (ViT) in MAE and extracts deep features $f_{\rm Deep}$ of masked patches. Its DNN predictor predicts SAR features of unseen patches. The target encoder uses the gradient-by-ratio method to map SAR images from pixel to feature space, thus extracting the target shapes and avoiding speckle noise interference in SAR images. Local patches and multi-scale features are designed for multi-scale small targets in remote sensing. For downstream datasets, we fine-tune the DNN encoder with a task layer using labeled training data. DNN encoder weights are frozen or fine-tuned using different tuning settings. Other model module weights are dropped. (c) The trained model is used to predict the test result.
The proposed Joint-Embedding Predictive Architecture for SAR target recognition (SAR-JEPA) is illustrated in Figure 2. Our overall objective is as follows: given local masked patches, predict feature representations in the SAR domain, w.r.t, the model can achieve SAR image interpretation by recognizing the contextual relationship between the target parts and the surrounding area.
As shown in Figure 2, the input is the single-channel SAR magnitude image since most target datasets are magnitude pictures. We randomly select different local patches and add mask tokens according to LoMaR (Chen et al., 2022a). Vision Transformers (ViT) (Dosovitskiy et al., 2020) in MAE are used as the DNN-encoder and DNN-predictor. Each local masked patch is fed into the MIM structure to predict the target feature of mask tokens. Importantly, the special SAR coherent imaging mechanism requires predicting high-quality target features instead of noisy pixels. SAR coherent imaging results in multiplicative speckle noise in the magnitude image, and we use the gradient-by-ratio (Dellinger et al., 2014; Song et al., 2016) to suppress noise and extract the targets’ edge and shape. Moreover, we employ a multi-scale and local mask setting for various targets in the pre-training set.
@@ 3.1 Local Masked Patches
We aim to learn the contextual relationship of various targets. However, such relationships normally remain in small areas due to the large scale and small target in remote sensing imagery. Therefore, we perform masking in the local patches following LoMaR. Although LoMaR is designed to improve computational efficiency, we found this local setting is suited for small targets in remote sensing. According to the LoMaR, mask tokens and relative positional encoding (Wu et al., 2021) are added to the encoder input and MAE structure to enhance the feature representation. These are important for LoMaR to avoid feature collapse.
@@ 3.2 Target Encoder
Due to the SAR imaging mechanism (Moreira et al., 2013), the images’ amplitude values contain multiplicative speckle noise. The speckle noise is related to the sensor resolution and is caused by the coherent superposition of different scatterers within a scattering cell. Therefore, reconstructing image pixels for SAR images receives noise interference and fails to learn a high-quality feature representation. This problem of speckle noise can be solved by filtering and feature extraction. However, filtering has a conflict between noise reduction and image details, and we use feature extraction. Some studies have proposed different types of target features in computer vision, mainly divided into traditional manual features (Wei et al., 2022) and deep learning features (Assran et al., 2023). However, these methods are difficult to migrate to SAR target recognition directly, and we combine the well-established local description feature methods (Dellinger et al., 2014; Song et al., 2016; Dong et al., 2020) in the SAR domain.
Gradient computation. The differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in the SAR image, and previous studies (Touzi et al., 1988; Bovik, 1988) have shown that the computing ratio is more suitable for multiplicative noise. Here, we use the simplest ratio of average (ROA) (Touzi et al., 1988) so as not to blur image details:
@@
where $R_{i}$ denotes the average ratio at different directions and $M_{1}(i)$ and $M_{2}(i)$ are the area averages on opposite sides of the current pixel along direction $i$. $i=1$ is the horizontal direction and $i=3$ means the vertical direction. As shown in Figure 2, the area averages $M$ can be computed from the input image with four fixed convolution kernels [^3] We added 1e-2 to the input image to prevent zero values.. Each convolution kernel extracts the sum of the region on one side of the pixel point, which is averaged to obtain $M$. By varying the size of the convolution kernel, the multi-scale features $f_{\rm SAR}$ can be obtained.
Then, the use of logarithms can solve the vertical gradient calculation (Dellinger et al., 2014), and the horizontal and vertical gradients are defined as:
@@
where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient. The gradient magnitude is $G_{\text{m}}=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}}$. We use gradient magnitude to construct multi-scale gradient features $f_{\rm SAR}$.
Multi-Scale Feature. SAR-HOG (Song et al., 2016) has discussed the kernel sizes for SAR vehicle target recognition and given the best single size for the MSTAR vehicle dataset. However, a single-size kernel cannot accommodate various targets when the dataset is extended to different targets and senors. Due to the dynamic range of contextual information for various targets in remote sensing (Li et al., 2023c), a multi-scale feature is constructed with convolutional kernels of different sizes in Figure 1.
We set the kernel size $r$ equal to 5, 7, 13, and 17, and the whole convolution kernel is a square with an odd $12r+$ 1. Though different kernel sizes, we have four gradient features $G_{\rm m1}$, $G_{\rm m2}$, $G_{\rm m3}$, and $G_{\rm m4}$, and the multi-scale feature $f_{\rm SAR}={\rm cat}(G_{\rm m1},G_{\rm m2},G_{\rm m3},G_{\rm m4})$ is a concatenat of four gradient features in the feature channel. As shown in Figure 1, small-scale convolution kernels provide a finer extraction of small target contours, while large-scale convolution kernels focus on large targets and scene edges and noise suppression.
@@ 3.3 Implementation
Given SAR images, we randomly sample several square windows at different spatial locations as local patches and mask each patch with mask tokens at a fixed percentage. Then, each local masked patch’s visible and unseen parts are fed into the MIM structure, whose encoder and predictor are applied with learnable relative positional encoding in the self-attention layer. Meanwhile, the whole SAR image is provided to the target encoder. The average values in four directions are obtained after multi-scale convolution in four directions. Then, the multi-scale SAR features are obtained after the ratio, logarithm, and magnitude operations. The loss function computes the mean squared error in the feature space between the DNN features $f_{\rm Deep}$ and SAR multi-scale features $f_{\rm SAR}$ for masked patches. In addition to the major improvement of the target encoder, there are two other minor.
@@
Decoder design. While some work (Chen et al., 2022a; Wei et al., 2022) has shown that it is feasible to use a simple linear head as a decoder or predictor for self-supervised learning in RGB images, we find that a ViT with eight-layers in the MAE is less disturbed by SAR image noise than a linear layer, allowing the encoder to extract deep semantic features of targets.
@@ 4 Experiments
This section analyzed the proposed architecture. We first described the pre-training and few-shot classification experimental setting in Sec. 4.1. Then, we discussed the effectiveness of the different modules in our architecture in Sec. 4.2. And the proposed SAR-JEPA was compared with other methods in Sec. 4.3 and in Sec. 4.4. In the end, Sec. 4.5 discussed the scaling experiment and SAR-JEPA showed improved performance with increasing unlabeled data.
@@ 4.1 Dataset and Experimental Settings
Table 1: Description of SAR datasets used for pre-training and downstream tasks. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Large SAR imagery in the detection contains more target and scene types than the annotation.
@@
MSTAR 5,216 10 1 0.3 X Single Fine-grained vehicle classification dataset
FUSAR-Ship 9,830 10 $\geq$5 1.1$\sim$1.7 C Double Fine-grained ship classification dataset
SAR-ACD 2,537 6 3 1 C Single Fine-grained aircraft classification dataset
Figure 3: Datasets for the pre-training and downstream tasks. Pre-training contain different targets, scenes, and sensors from MSAR, SAR-Ship-Dataset, SARSim, and SAMPLE. MSAR is the satellite-based dataset of ground and sea targets; SAR-Ship-Dataset is the satellite-based dataset of sea targets; SARSim is the multi-angle simulation dataset of vehicle targets; and SAMPLE is the simulated and measured dataset of vehicle targets. We use MSTAR, FUSAR-Ship, and SAR-ACD datasets to evaluate the performance for recognizing different targets. MSTAR is a fine-grained vehicle dataset; FUSAR-Ship is a sea targets dataset; SAR-ACD is a fine-grained aircraft dataset.
Pre-training. Pre-training datasets are designed to ensure rich information about targets, scenes, and sensors. Therefore, we chose four datasets to build a pre-training dataset for SSL. It covers popular targets (vehicles, ships, aircraft, etc.), complex backgrounds (ground and ocean), and various sensors (satellite-borne, airplane-borne, and simulation) in Table 1 and Figure  3.
@@
Experimental settings. TTable 2 gives the default setting. The pre-training is applied on 4 NVIDIA V100 GPUs with 200 epochs and 300 batch sizes, and the backbone model is ViT-Base (ViT-B). Other specific hyperparameters of each method use the default settings from their papers and codes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness, and all methods use the same data augmentation. Since our goal is not to get better performance by tuning downstream hyperparameters, all models use the same training settings in downstream classification tasks. The few-shot learning setting is based on the Dassl toolbox (Zhou et al., 2023a) and is averaged over 10 random experiments.
@@
4.2 Ablation Study
Here, we analyze three key points of the proposed method. The first compares local (LoMaR) and global (MAE) mask methods and explains why we apply LoMaR with its modified ViT and MAE’s deep decoder as the backbone. Then, we discuss different target features (LPF, HOG, SAR-HOG, and ours) on the same backbone to show the need for suppressing noise and target features. Finally, we discuss the role of MIM and PGCA in our architecture.
Table 3: Discussion of the MIM framework for SAR-JEPA. MAE performs global masks, and LoMaR performs local masks. LoMaR-SAR is our modified LoMaR with decoder depth, and we use it as our SAR-JEPA backbone. The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
fine-tuning
Method MSTAR FUSAR-Ship SAR-ACD
MAE (He et al., 2022) 50.6 61.3 69.5 71.7 75.4 78.2 51.6 57.0 69.5
LoMaR (Chen et al., 2022a) 45.6 62.9 77.0 75.9 80.2 82.7 51.2 54.4 67.4
LoMaR-SAR* 55.3 69.8 84.5 78.6 82.0 84.8 51.7 60.1 71.7
linear probing
Method MSTAR FUSAR-Ship SAR-ACD
MAE (He et al., 2022) 53.7 59.2 63.7 74.2 78.1 80.3 49.9 56.9 62.7
LoMaR (Chen et al., 2022a) 39.3 48.9 55.6 70.5 77.9 80.7 47.8 53.4 59.1
LoMaR-SAR* 57.7 64.4 72.2 78.0 81.0 84.0 54.4 58.9 64.4
    We replace the linear decoder of LoMaR with an eight-layer decoder in MAE to reduce the effect of noise in the reconstructed pixels by increasing the decoder depth. This improvement inspired us to resolve thoroughly SAR image noise with high-quality target features.
Table 4: Results of different target features for SAR-JEPA. LPF: low pass filter. HOG: histogram of oriented gradient. GR: gradient-by-ratio. The baseline is pixel value, which is corrponds to LoMaR-SAR in Table 3. LPF uses the default filter hyperparameters in (Liu et al., 2023). SAR-HOG and two GR method, whose kernel size can be adjusted, use the same multi-scale setting. $\rm{GR_{lin}}$: gradient-by-ratio with linear kernel. $\rm{GR_{Gau}}$: gradient-by-ratio with Gaussian kernel. The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
fine-tuning
Target feature MSTAR FUSAR-Ship SAR-ACD
Pixel 55.3 69.8 84.5 78.6 82.0 84.8 51.7 60.1 71.7
LPF (Liu et al., 2023) 57.7 72.7 86.5 78.4 81.9 84.9 51.9 61.2 72.8
HOG (Wei et al., 2022) 30.2 32.4 40.3 58.1 63.3 66.6 51.7 57.9 64.7
SAR-HOG (Song et al., 2016) 64.4 76.4 90.5 76.7 80.8 85.3 51.4 60.0 73.4
$\rm{GR_{Gau}}$* (Dellinger et al., 2014) 62.8 78.5 88.2 80.6 82.8 86.0 55.3 62.9 73.0
$\rm{GR_{lin}}$ (Song et al., 2016) 70.3 82.1 91.6 81.3 83.3 85.8 54.8 62.6 75.5
linear probing
Target feature MSTAR FUSAR-Ship SAR-ACD
Pixel 57.7 64.4 72.2 78.0 81.0 84.0 54.4 58.9 64.4
LPF (Liu et al., 2023) 56.1 62.6 68.1 78.7 82.0 84.6 54.9 59.5 65.2
HOG (Wei et al., 2022) 33.3 37.8 41.6 52.6 58.3 61.5 49.2 53.1 56.7
SAR-HOG (Song et al., 2016) 58.9 68.1 77.1 79.3 81.8 85.0 54.6 59.3 66.5
$\rm{GR_{Gau}}$* (Dellinger et al., 2014) 62.8 72.3 77.6 82.0 82.9 85.7 54.7 60.9 66.6
$\rm{GR_{lin}}$ (Song et al., 2016) 67.7 75.1 81.6 80.6 82.7 85.2 56.8 60.4 66.8
    We finally chose the linear kernel $\rm{GR_{lin}}$ for SAR-JEPA because it needed a lot of time to fine-tune the Gaussian kernel $\rm{GR_{Gau}}$ parameters in pre-training. And we set the Gaussian standard deviation from $\sigma=0.3*((2r+1-1)*0.5-1)+0.8$. It means the standard deviation is related to the kernel size instead of a fixed parameter in Dellinger et al. (2014).
Figure 4: Multi-scale kernel settings for gradient-by-ratio $\rm{GR_{lin}}$. Here, the scale 1/2/3/4 is with $r$ equal to 5/7/13/17, and the multi-scale concat all four scales in feature channel. Multi-scale is more suitable than single scale due to various targets in remote sensing images.
Table 5: Ablation study of our SAR-JEPA, which is a combination of two architectures (MIM and PGCA). The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
fine-tuning
Method MSTAR FUSAR-Ship SAR-ACD
MIM* 55.3 69.8 84.5 78.6 82.0 84.8 51.7 60.1 71.7
PGCA** 56.7 71.4 88.1 76.4 80.4 83.8 51.8 60.0 73.7
SAR-JEPA 70.3 82.1 91.6 81.3 83.3 85.8 54.8 62.6 75.5
linear probing
Method MSTAR FUSAR-Ship SAR-ACD
MIM* 57.7 64.4 72.2 78.0 81.0 84.0 54.4 58.9 64.4
PGCA** 61.0 68.0 74.7 77.4 81.6 84.2 52.7 58.3 66.1
SAR-JEPA 67.7 75.1 81.6 80.6 82.7 85.2 56.8 60.4 66.8
    MIM is refers to LoMaR-SAR in Table 3, which is also the MIM module of our SAR-JEPA.
    Due to PGCA (Datcu et al., 2023) using complex SAR image’s features as target signal, we follow its ideas and use our target encoder. Here, PGCA is the contrastive learning between deep features and our multi-scale gradient features $\rm{GR_{lin}}$ to test the effectiveness of the target feature as a guided signal, w.r.t, it removes the MIM task in SAR-JEPA.
LoMaR vs. MAE. Since most targets are small objects in remote sensing images, we prefer masking in local areas to learn the contextual relationships near the target better. Compared to MAE, which masks over the whole image, LoMaR randomly selects multiple local squares for masking. In addition to this, LoMaR uses relative positional encoding, input with mask tokens, and linear decoder. As LoMaR emphasizes, the first two are indispensable to prevent feature collapse. When we remove any of them, pre-training in SAR images with LoMaR and LoMaR-SAR have performance degradation and feature collapse.
LoMaR uses a linear decoder to speed up the computation. However, the linear layer is unsuitable for SAR images because SAR image noise is easier than a deep decoder to affect the encoder features through the loss function. In our early experiments, LoMaR performed better than MAE on small batches of high-quality pre-training data. When we extended the pre-training set to different quality datasets, LoMaR’s performance became unstable. This phenomenon inspired us to make improvements to the decoder and target features. As shown in Table 3, LoMaR-SAR, which replaces the linear decoder of LoMaR with an eight-layer decoder in MAE [^4] However, deeper decoders can be detrimental to encoder performance. For example, the twelve-layer decoder is worse than the eight-layer in the MAE paper., has the best result. Therefore, we use LoMaR-SAR as the MIM module for SAR-JEPA. The target encoder for SAR-JEPA is discussed below. It needs unique designs for SAR images and is at the heart of our architecture.
Target feature selection. In Table 4, we compare the results of pixel value, Low Pass Filter (LPF) (Liu et al., 2023), SAR-HOG (Song et al., 2016), and Gradient-by-Ratio (GR) (Dellinger et al., 2014). All the changes are made only in the target features, while other settings remain unchanged. The Pixel in Table 4 is LoMaR-SAR as baseline. The LPF can mitigate noise interference, but we observe a slight drop of LPF compared to Pixel on the MSTAR’s linear probing result. Therefore, inspired by the HOG feature in MaskFeat, we prefer to obtain high-quality target features with suppressing noise.
Although HOG features have shown effectiveness in FG-MAE for SAR scene-level SSL, we found that the HOG features are unsuitable for target-level SSL in Table 4. This result is because speckle noise can cause the gradient computation in small target regions to have many noise points, greatly affecting the quality of target features. Inspired by the physics-guided contrastive architecture using SAR physics features, we investigate the HOG feature variant in the SAR domain, i.e., SAR-HOG. The core of SAR-HOG is the change in the gradient calculation from differential gradient to ratio gradient. As shown in Table 4, we observe a significant improvement compared to HOG due to this modification. However, the sampling method of HOG may lead to the loss of some small target scattering points, such as the discrete points of an aircraft or a small boat. Therefore, we focus on the gradient information, i.e., target edges. Moreover, basic information, such as the shape and edges of an object, is a generic property (Liu et al., 2022; Cheng et al., 2022; Shi et al., 2020).
We consider two versions of GR in SAR. $\rm{GR_{lin}}$ is computing gradient with a linear kerenal (Song et al., 2016) whose $M$ value equals 1. $\rm{GR_{Gau}}$ is computing gradient with a Gaussian kernel (Dellinger et al., 2014) whose $M$ value obeys a Gaussian distribution. Table 4 shows that the two calculation methods have their advantages for different target categories and have similar performance in most cases. The Gaussian kernel can act as a filter, but the parameters need to be adjusted so as not to blur the target edges overly. Because a pre-trained model needs tens of hours to several days, we prefer to use simple and effective methods with scaling data and parameters. Therefore, we use a linear kernel without spending too much time tuning the parameters of the Gaussian kernel at different scales.
Another important thing is that we propose a multi-scale setting instead of single-scale features for various targets included in the pre-training set. Multi-scale can improve the feature representation for various targets in remote sensing. Figure 4 gives the result of different kernel settings. Scale settings have various effects: smaller scales can better extract small target edges, while larger scales are more obvious for noise suppression. Hence, learning target features at different scales is more advantageous than a single scale.
SAR-JEPA is a combination of MIM and PGCA ideas as shown in Figure 1. Specifically, we designed LoMaR-SAR and multi-scale gradient features as a concrete implementation of the MIM and target encoder module. Table 5 demonstrates the equal importance and mutual reinforcement of the two modules. These focus on the pretext task and signal quality of SSL driven by the data. MIM can provide PGCA with a challenging pretext task of learning features with contextual information in the image. PGCA can improve the quality of features by introducing domain knowledge into the embedding of the target encoder. Combining MIM and PGCA, SAR-JEPA learns high-quality contextual features for low-quality noisy SAR data.